OpenStudy (brittany16):

Solve for the variable. 37 = -5 + y

OpenStudy (anonymous):

y=42

jimthompson5910 (jim_thompson5910):

37 = -5 + y 37 + 5 = -5 + y + 5 42 = y y = 42